OpenStudy (mackenzie2013):

The slope of the graph of y = 5x is ?

OpenStudy (anonymous):

5, x=slope

OpenStudy (anonymous):

the standard form for a linear equation is: y=mx+b, mx=slope, and b=y-intercept.

OpenStudy (anonymous):

Excuse me? I don't see -2 anywhere.

OpenStudy (mackenzie2013):

Alright (0,0) (1,5)

OpenStudy (anonymous):

5-0/1-0=5/1. Slope: 5 :)

OpenStudy (mackenzie2013):

Thank you!

OpenStudy (anonymous):

No problem! Anymore?